# Rotationally invariant dynamical lattice regulators for Euclidean quantum field theories

### v1

@@ 3.3 Gauge invariance and lattice symmetries
By construction the prototype action $S_{\mathrm{fields}}^{\circ}[x,U,\Phi]$ is a sum of gauge–invariant local terms built from traces of Wilson loops and gauge–covariant combinations of $\Phi$ and its lattice covariant derivatives. For example, the gauge part of the action may be written as
$S_{g}^{\circ}[x,U]=\sum_{n,\mu<\nu}w^{\mu\nu}(x;n)\,\Re\mathrm{tr}\bigl(\mathbf{1}-U_{\mu\nu}(n)\bigr),$ (3.46)
with local geometry–dependent weights $w_{\mu\nu}(x;n)$ and plaquette variables $U_{\mu\nu}(n)$ built from the links $U_{\mu}(n)$ in the usual way. Under (3.43) the plaquettes transform by conjugation and the traces are invariant. Likewise, the matter part of $S_{\mathrm{fields}}^{\circ}$ is built from gauge–covariant bilinears such as $\bar{\Phi}(n)\,\Phi(n)$ and $\bar{\Phi}(n)\,D_{\mu}\Phi(n)$, which are invariant under (3.44). Since $x$ is gauge neutral and appears only through gauge–invariant weights such as $w_{\mu\nu}(x;n)$, we have
@@
Lattice translations and hypercubic rotations. In addition to gauge transformations the theory is invariant under the usual discrete symmetries of the abstract hypercubic lattice $\Lambda$. Let $k\in\mathbb{Z}^{d}$ be a lattice translation vector. The translation $\tau_{k}$ acts on the fields by
$(\tau_{k}x)(n):=x(n+k),\qquad(\tau_{k}U)_{\mu}(n):=U_{\mu}(n+k),\qquad(\tau_{k}\Phi)(n):=\Phi(n+k).$ (3.48)
Similarly, let $R_{\mathrm{lat}}$ be an element of the hypercubic group $H_{d}\subset\mathrm{GL}(d,\mathbb{Z})$, i.e. an orthogonal integer matrix representing a permutation and/or sign flip of the lattice axes. Then the induced hypercubic transformation acts by
$(R_{\mathrm{lat}}x)(n):=x(R_{\mathrm{lat}}^{-1}n),\qquad(R_{\mathrm{lat}}U)_{\mu}(n):=U_{\rho}\bigl(R_{\mathrm{lat}}^{-1}n\bigr),\qquad(R_{\mathrm{lat}}\Phi)(n):=\Phi\bigl(R_{\mathrm{lat}}^{-1}n\bigr),$ (3.49)
with $\rho$ the image of the direction $\mu$ under $R_{\mathrm{lat}}$. Because the action is a sum of identical local contributions over sites, links, and plaquettes of the abstract lattice, and the geometry weights $g(n)$ and $V(n)$ are evaluated at the corresponding transformed sites, the full action and measure are invariant under these discrete lattice symmetries.
@@
where each $f_{n}$ depends only on the restriction of $(x,U,\Phi)$ to a finite neighbourhood of $n$, with a radius (in lattice distance) that is independent of $n$ and of the lattice size. In particular, the geometry action $S_{x}[x]$ in (2.18) and the field action $S_{\mathrm{fields}}[x,U,\Phi]$ are local in this sense: the local density at site $n$ depends on the geometry only through the frame $E(n)$ (or a bounded stencil of neighbouring frames) and on the gauge/matter fields only through fields on a fixed finite stencil of sites and links around $n$.
@@ 3.4 Viewpoint as an extended lattice field theory
and the geometry–dependent quantities $e_{\mu}(n)$, $g_{\mu\nu}(n)$ and $V(n)$ are just local composite fields built from the $x_{\mu}(n)$ (or $X_{\mu}(n)$) on a finite stencil. The action $S[x,U,\Phi]$ defined in (3.1) is then a sum of strictly local terms on this extended field space, and the path integral (3.2) is an ordinary Euclidean lattice path integral with local interactions.
In this extended viewpoint the main structural properties of the regulator follow from standard lattice arguments applied to the enlarged field content. Locality and gauge invariance are manifest at the level of the abstract graph, as discussed in Section 3.3. Reflection positivity is proved by choosing a reflection of the time index $n_{1}$ on $\Lambda$ and checking the usual Osterwalder–Schrader inequalities for the extended field multiplet $(x,U,\Phi)$; the detailed construction of the reflection, the splitting of the lattice into time slices, and the verification of the OS axioms are carried out in Section 4. In particular, no continuum notion of “time” or preferred coordinate direction in $\mathbb{R}^{d}$ is required at the microscopic level: all reflection operations are defined purely on the index lattice, and the embedding variables $x_{\mu}(n)$ enter only through local, reflection–invariant combinations. This perspective makes clear that the dynamical–lattice regulator fits squarely within the usual constructive framework for Euclidean lattice field theory, with the dynamical geometry fields providing an $\mathrm{E}(d)$–invariant way of dressing local couplings without sacrificing locality or reflection positivity.
@@ 4 Reflection positivity and continuum universality
In this section we first establish Osterwalder–Schrader (OS) reflection positivity for the dynamical–lattice regulator, and briefly recall the associated Hilbert–space reconstruction. We then discuss the Symanzik effective–action description of the continuum limit and explain why, under mild assumptions on the geometry sector, the dynamical–lattice theory lies in the same universality class as the corresponding theory on a static hypercubic lattice. As a concrete check we sketch the one–loop running of the quartic coupling in scalar $\phi^{4}$ theory.
@@ 4.1 Reflection positivity for the bosonic sector
where $r\in\mathrm{O}(d)$ is the target–space reflection $r(v_{1},v_{2},\dots,v_{d})=(-v_{1},v_{2},\dots,v_{d})$, the reflection $\theta\mu$ on the direction $\mu$ is as in (3.4), and the overline denotes complex conjugation.
@@
$\begin{split}e_{\mu}(\Theta x;n)&=(\Theta x)(n+\hat{\mu})-(\Theta x)(n)=rx(\theta(n+\hat{\mu})-rx(\theta n)\\ &=r\bigl(x(\theta n+\widehat{\theta\mu})-x(\theta n)\bigr)=r\,e_{\theta\mu}(x;\theta n).\end{split}$ (4.4)
@@
We now state the reflection–positivity criterion for the coupled DLR measure. In the concrete examples of Section 3.1, the total action $S=S_{x}+S_{\mathrm{fields}}$ is local on the abstract lattice and is $\Theta$–invariant by construction. In the theorem we abstract these features and assume only locality, $\Theta$–invariance, and finite interaction range in the index–time direction.
@@
where the overline denotes complex conjugation (in particular, $\Theta$ is antilinear). With the above support convention, $\Theta$ maps $\mathcal{A}_{+}$ into the corresponding algebra on $\Lambda_{-}$.
Here and below, for functionals $F,G$ on configuration space we write $(F\cdot G)(\omega):=F(\omega)\,G(\omega)$ for their pointwise product, so that $\langle\Theta F\cdot F\rangle=\int(\Theta F)(\omega)\,F(\omega)\,d\mu(\omega)$.
@@
$S[\Theta x,\Theta U,\Theta\Phi]\;=\;S_{\mathrm{tot}}[x,U,\Phi]\qquad\text{for all }x\in\mathcal{X}_{\mathrm{adm}},$ (4.11)
and that the interaction range in the time direction is finite. Then we have
@@
We prove (4.12) by the standard factorization argument for link reflection, specialized to the nearest–neighbour stencils used in § 3.1. To this end we isolate the finite “reflection slab”
$\Sigma=\{n\in\Lambda:\ n_{1}\in\{0,1\}\}.$ (4.13)
We decompose a configuration as $\omega=(\omega_{-},\omega_{0},\omega_{+})$ by declaring: (i) $\omega_{0}$ consists of all site variables $(x(n),\Phi(n))$ with $n\in\Sigma$, together with all link variables $U_{\mu}(n)$ whose link $\{n,n+\hat{\mu}\}$ meets $\Sigma$ (that is, $n\in\Sigma$ or $n+\hat{\mu}\in\Sigma$); (ii) $\omega_{+}$ consists of the remaining variables with support in the strict future $\{n_{1}\geq 2\}$ (sites) and links with both endpoints in $\{n_{1}\geq 2\}$; and (iii) $\omega_{-}$ consists of the remaining variables with support in the strict past $\{n_{1}\leq-1\}$ (sites) and links with both endpoints in $\{n_{1}\leq-1\}$.
For the hard constraint we use the convention that $S(\omega)=+\infty$ whenever $x\notin\mathcal{X}_{\mathrm{adm}}$, so that all locality statements below include the constraint. In the concrete actions of § 3.1, every local interaction term has time-range one, hence depends on variables from at most two adjacent time slices. Moreover, using the action–averaged prescription (each prototype term is paired with its reflected copy), the $\Theta$–invariant part supported in the slab is completely explicit:
(Geometry) the slab contribution consists of the geometry terms whose stencil uses only sites in $\Sigma$ (in particular, all terms involving the time edge between the slices 0 and $1$), together with their reflected partners.
(Scalars/fermions) the slab contribution consists of all on–site potential/mass terms on the slices $n_{1}=0,1$, and the nearest–neighbour kinetic couplings across the reflected time link between $n_{1}=0$ and $n_{1}=1$ (again including both members of each $(\text{prototype},\Theta\text{-copy})$ pair). All other kinetic terms couple $\omega_{0}$ only to $\omega_{+}$ (for sites in the slice $n_{1}=1$) or only to $\omega_{-}$ (for sites in the slice $n_{1}=0$).
(Gauge fields) the slab contribution consists of all Wilson loop terms built from links meeting the slices $n_{1}=0,1$ only: in particular, all purely spatial plaquette terms in the slices $n_{1}=0$ and $n_{1}=1$, and all mixed (time-space) plaquette terms that use the reflected time link between the slices 0 and $1$, together with their reflected partners. Plaquette terms based at $n_{1}=1$ that extend to the slice $n_{1}=2$ are not in the slab and belong to the positive part below.
With this bookkeeping the total action admits a decomposition
$S(\omega)=S_{+}(\omega_{0},\omega_{+})+S_{0}(\omega_{0})+S_{-}(\omega_{0},\omega_{-}),\qquad S_{-}=\Theta S_{+},$ (4.14)
where $S_{0}$ is the sum of the slab terms described above (hence depends only on $\omega_{0}$) and is $\Theta$–invariant, while $S_{+}$ (respectively $S_{-}$) collects the remaining local terms supported in the future (respectively past), together with those terms in the slice $n_{1}=1$ (respectively $n_{1}=0$) that couple $\omega_{0}$ to $\omega_{+}$ (respectively to $\omega_{-}$). In our concrete construction one also has $S_{0}=\Theta S_{0}$, but we do not need this separately.
Now let $F\in\mathcal{A}_{+}$, so $F$ depends only on the variables in $\omega_{0}$ and $\omega_{+}$. Define
@@
where $d\omega_{+}$ denotes the product measure over the variables comprising $\omega_{+}$. Using $\Theta$–invariance of the domain and product measures together with (4.14) and $(\Theta F)(\omega)=\overline{F(\Theta\omega)}$, one obtains
@@
which is (4.12). For actions with larger time stencils one enlarges $\Sigma$ by the corresponding finite number of adjacent time slices so that the same bookkeeping applies. ∎
@@
  1. (i)
    Theorem 4.3 is the correct positivity statement for the DLR: the embedding $x$ is part of the Euclidean field content, and the reconstructed Hilbert space consists of wavefunctionals of both geometry and matter boundary data. In particular one should not expect reflection positivity to hold at fixed, asymmetric $x$.
  2. (ii)
    The symmetrized admissibility conditions in Definition 2.1 ensure $\Theta(\mathcal{X}_{\rm adm})=\mathcal{X}_{\rm adm}$, so the change of variables $\omega\mapsto\Theta\omega$ used in the factorization argument is legitimate.
  3. (iii)
    Only one reflection (here: link reflection) is needed for OS reconstruction; site and link reflections, when both available, differ by a one-step lattice translation.
  4. (iv)
    For the models of §3.1 the hypotheses of Theorem 4.3 are satisfied: ultra-local potentials are manifestly $\Theta$-invariant, and the kinetic/Wilson terms are built from reflection-covariant geometric data together with the standard link-reflection rule for $U$.
4.2 Wilson fermions and reflection positivity
We extend the coupled DLR measure to include fermions by introducing Grassmann-valued lattice spinors. Work at finite volume on the cylinder $\Lambda_{T,L}$ as in §4.1. At each site $n\in\Lambda_{T,L}$ let $\psi(n)$ and $\bar{\psi}(n)$ be independent Grassmann generators, carrying spin and internal indices and transforming in a unitary representation of $G$. The fermionic integration is the Berezin product measure
$D\bar{\psi}\,D\psi=\prod_{n\in\Lambda_{T,L}}\prod_{\alpha}d\bar{\psi}_{\alpha}(n)\,d\psi_{\alpha}(n),$

### v2

@@ 3.3 Gauge invariance and lattice symmetries
By construction the prototype action $S_{\mathrm{fields}}^{\circ}[x,U,\Phi]$ is a sum of gauge invariant local terms built from traces of Wilson loops and gauge covariant combinations of $\Phi$ and its lattice covariant derivatives. For example, the gauge part of the action may be written as
$S_{g}^{\circ}[x,U]=\sum_{n,\mu<\nu}w^{\mu\nu}(x;n)\,\Re\mathrm{tr}\bigl(I-U_{\mu\nu}(n)\bigr),$ (3.49)
with local geometry–dependent weights $w_{\mu\nu}(x;n)$ and plaquette variables $U_{\mu\nu}(n)$ built from the links $U_{\mu}(n)$ in the usual way. Under (3.46) the plaquettes transform by conjugation and the traces are invariant. Likewise, the matter part of $S_{\mathrm{fields}}^{\circ}$ is built from gauge covariant bilinears such as $\bar{\Phi}(n)\,\Phi(n)$ and $\bar{\Phi}(n)\,D_{\mu}\Phi(n)$, which are invariant under (3.47). Since $x$ is gauge neutral and appears only through gauge invariant weights such as $w_{\mu\nu}(x;n)$, we have
@@
Lattice translations and hypercubic symmetries. In addition to gauge transformations, the theory is invariant under lattice translations and under those hypercubic relabellings of the abstract lattice that preserve the admissible domain $\mathcal{X}_{\mathrm{adm}}(a)$. Let $\tau_{k}:\Lambda\to\Lambda$ be the translation $\tau_{k}(n)=n+k$, and let
$\mathrm{H}(d)=\{R\in\mathrm{GL}(d,\mathbb{Z}):R^{\mathsf{T}}R=I\},\qquad\mathrm{H}^{+}(d)=\{R\in\mathrm{H}(d):\det R=+1\}.$ (3.51)
We let these maps act on fields by pullback:
$(\tau_{k}x)(n):=x(n-k),\qquad(\tau_{k}\Phi)(n):=\Phi(n-k),\qquad(\tau_{k}U)(\ell):=U(\tau_{k}^{-1}\ell),$ (3.49)
and for $R_{\rm lat}\in\mathrm{H}(d)$,
$(R_{\rm lat}x)(n):=x(R_{\rm lat}^{-1}n),\qquad(R_{\rm lat}\Phi)(n):=\Phi(R_{\rm lat}^{-1}n),\qquad(R_{\rm lat}U)(\ell):=U(R_{\rm lat}^{-1}\ell),$ (3.50)
[…]
With the basic definition of $\mathcal{X}_{\mathrm{adm}}(a)$ in terms of a distinguished corner frame and a selected orientation component, we therefore only claim invariance under the corresponding subgroup of $\mathrm{H}^{+}(d)$ that fixes that corner choice. If instead one imposes the strengthened all–corner oriented–volume condition of Remark 2.4(A3), then $\mathcal{X}_{\mathrm{adm}}(a)$ is invariant under the full orientation–preserving hypercubic group $\mathrm{H}^{+}(d)$. Permutations with $\det=-1$ exchange the two orientation components and hence would not be included.
@@
where each $f_{n}$ depends only on the restriction of $(x,U,\Phi)$ to a finite neighbourhood of $n$, with a radius (in lattice distance) that is independent of $n$ and of the lattice size. In particular, the geometry action $S_{x}[x]$ in (2.20) and the field action $S_{\mathrm{fields}}[x,U,\Phi]$ are local in this sense: the local density at site $n$ depends on the geometry only through the frame $E(n)$, or a bounded stencil of neighbouring frames, and on the gauge/matter fields only through fields on a fixed finite stencil of sites and links around $n$.
@@ 3.4 Viewpoint as an extended lattice field theory
and the geometry–dependent quantities $e_{\mu}(n)$, $g_{\mu\nu}(n)$ and $V(n)$ are just local composite fields built from the $x_{\mu}(n)$ on a finite stencil. The action $S[x,U,\Phi]$ defined in (3.1) is then a sum of strictly local terms on this extended field space, and the path integral (3.2) is an ordinary Euclidean lattice path integral with local interactions.
In this extended viewpoint the main structural properties of the regulator follow from standard lattice arguments applied to the enlarged field content. Locality and gauge invariance are manifest at the level of the abstract graph, as discussed in Section 3.3. Reflection positivity is proved by choosing a reflection of the time index $n_{1}$ on $\Lambda$ and checking the usual Osterwalder–Schrader inequalities for the extended field multiplet $(x,U,\Phi)$; the detailed construction of the reflection, the splitting of the lattice into time slices, and the verification of the OS axioms are carried out in Section 4. In particular, no continuum notion of “time” or preferred coordinate direction in $\mathbb{R}^{d}$ is required at the microscopic level: all reflection operations are defined purely on the index lattice, and the embedding variables $x_{\mu}(n)$ enter only through local, reflection invariant combinations. This perspective makes clear that the dynamical–lattice regulator fits squarely within the usual constructive framework for Euclidean lattice field theory, with the dynamical geometry fields providing an $\mathrm{E}(d)$–invariant way of dressing local couplings without sacrificing locality or reflection positivity.
@@ 4 Reflection positivity and continuum universality
In this section we first establish Osterwalder–Schrader (OS) reflection positivity for a large class of DLR, and briefly recall the associated Hilbert space reconstruction. We then discuss the Symanzik effective action description of the continuum limit and explain why, under mild assumptions on the geometry sector, the dynamical–lattice theory lies in the same universality class as the corresponding theory on a static hypercubic lattice. As a concrete check we sketch the one–loop running of the quartic coupling in scalar $\phi^{4}$ theory.
@@ 4.1 Reflection positivity for the bosonic sector
In this subsection we treat the bosonic matter sector, i.e., gauge fields coupled to complex scalar fields $\phi$; the fermionic sector is treated separately in the next subsection.
@@
where $r\in\mathrm{O}(d)$ is the target space reflection $r(v_{1},v_{2},\dots,v_{d})=(-v_{1},v_{2},\dots,v_{d})$, the reflection $\theta\mu$ on the direction $\mu$ is as in (3.4), and the overline denotes complex conjugation.
@@
$\begin{split}e_{\mu}(\Theta x;n)&=(\Theta x)(n+\hat{\mu})-(\Theta x)(n)=rx(\theta(n+\hat{\mu}))-rx(\theta n)\\ &=r\bigl(x(\theta n+\widehat{\theta\mu})-x(\theta n)\bigr)=r\,e_{\theta\mu}(x;\theta n).\end{split}$ (4.4)
@@
We state the reflection–positivity criterion for the coupled DLR measure. In the concrete examples of §3.1, the total action $S=S_{x}+S_{\mathrm{fields}}$ is local on the abstract lattice and is $\Theta$–invariant by construction (each prototype term is paired with its $\theta$–partner and then averaged), and the admissible set $\mathcal{X}_{\mathrm{adm}}(a)$ is $\Theta$–stable. In the theorem we abstract these features and assume only locality, $\Theta$–invariance, and finite interaction range in the index–time direction.
@@
where the overline denotes complex conjugation. In particular, $\Theta$ is antilinear. With the above support convention, $\Theta$ maps $\mathcal{A}_{+}$ into the corresponding algebra on $\Lambda_{-}$.
Here and below, for functionals $F,G$ on configuration space we write $(F\cdot G)(\omega)=F(\omega)\,G(\omega)$ for their pointwise product, so that $\langle\Theta F\cdot F\rangle=\int(\Theta F)(\omega)\,F(\omega)\,d\mu(\omega)$.
@@
$S[\Theta x,\Theta U,\Theta\phi]=S[x,U,\phi]\qquad\text{for all }x\in\mathcal{X}_{\mathrm{adm}}(a),$ (4.11)
and that the interaction has finite range in the index–time direction: there exists an integer $\ell\geq 1$ such that each local term in $S$ depends only on variables whose site time–indices $n_{1}$ lie in an interval of length at most $\ell$ (equivalently, the $n_{1}$–diameter of the stencil is $\leq\ell$). Then
@@
We use the standard link–reflection factorization argument. Fix $\ell$ as in the hypothesis and introduce the finite reflection slab
$\Sigma_{\ell}=\{\,n\in\Lambda:\ 1-\ell\leq n_{1}\leq\ell\,\}.$
For the nearest–neighbour/plaquette actions of §3.1 one has $\ell=1$, hence $\Sigma_{\ell}=\{n_{1}\in\{0,1\}\}$. Decompose a configuration as $\omega=(\omega_{-},\omega_{0},\omega_{+})$ by declaring: (i) $\omega_{0}$ consists of all site variables $(x(n),\phi(n))$ with $n\in\Sigma_{\ell}$, together with all link variables $U_{\mu}(n)$ whose link $\{n,n+\hat{\mu}\}$ meets $\Sigma_{\ell}$; (ii) $\omega_{+}$ consists of the remaining variables with support in the strict future $\{n_{1}\geq\ell+1\}$ (sites) and links with both endpoints in $\{n_{1}\geq\ell+1\}$; and (iii) $\omega_{-}$ consists of the remaining variables with support in the strict past $\{n_{1}\leq-\ell\}$ (sites) and links with both endpoints in $\{n_{1}\leq-\ell\}$.
By the finite time–range assumption, no local interaction term can simultaneously involve variables from the strict past $\{n_{1}\leq-\ell\}$ and the strict future $\{n_{1}\geq\ell+1\}$. Hence every local term belongs to exactly one of the following three classes: it is supported entirely in $\Sigma_{\ell}$; or it involves at least one variable from the strict future; or it involves at least one variable from the strict past. Therefore the total action decomposes as
$S(\omega)=S_{+}(\omega_{0},\omega_{+})+S_{0}(\omega_{0})+S_{-}(\omega_{0},\omega_{-}),$ (4.13)
where $S_{0}$ is the sum of all terms supported in $\Sigma_{\ell}$, and $S_{+}$ (resp. $S_{-}$) collects all remaining terms that involve at least one strict-future (resp. strict-past) variable; in particular $S_{+}$ depends only on $(\omega_{0},\omega_{+})$ and $S_{-}$ only on $(\omega_{0},\omega_{-})$. Using that $\mathcal{X}_{\mathrm{adm}}$ is $\Theta$–stable and that the product measures are invariant under the induced relabelings, together with $(\Theta F)(\omega)=\overline{F(\Theta\omega)}$, one obtains
$S_{-}(\omega_{0},\omega_{-})=(\Theta S_{+})(\omega_{0},\omega_{-}).$ (4.14)
Now let $F\in\mathcal{A}_{+}$, so $F$ depends only on the variables in $\Lambda_{+}$ and thus only on $(\omega_{0},\omega_{+})$. Define
@@
where $d\omega_{+}$ denotes the product measure over the variables comprising $\omega_{+}$. Using $\Theta$–invariance of the domain and product measures together with (4.13)–(4.14) and $(\Theta F)(\omega)=\overline{F(\Theta\omega)}$, we obtain
@@
which is (4.12). ∎
@@
Only one reflection plane is needed for OS reconstruction. Site and link reflections, when both are available, differ by a one–step lattice translation. Here we work with the link reflection between the slices $n_{1}=0$ and $n_{1}=1$.
4.2 Fermions and reflection positivity
We extend the coupled DLR measure to include fermions by introducing Grassmann-valued lattice spinors. Work at finite volume on the cylinder $\Lambda=\Lambda_{T,L}$ as in the preceding subsection, with the understanding that statements on the infinite lattice are interpreted by taking the limit $T,L\to\infty$ on local observables. At each site $n\in\Lambda$ let $\psi(n)$ and $\bar{\psi}(n)$ be independent Grassmann generators, carrying spin and internal indices and transforming in a unitary representation of $G$. The fermionic integration is the Berezin product measure
$D\bar{\psi}\,D\psi=\prod_{n\in\Lambda}\prod_{\alpha}d\bar{\psi}_{\alpha}(n)\,d\psi_{\alpha}(n),$